OpenStudy (anonymous):

Issac and Hailee drove to an amusement park that is 48 miles away. Their average speed was 53 miles per hour. How long was the total travel time to the nearest tenth of an hour?

OpenStudy (anonymous):

Please help.

OpenStudy (anonymous):

if they make 53 miles an hour, and the park is only 48 miles away, it took them less than an hour right?

OpenStudy (anonymous):

Yes, they did. Thanks again.

OpenStudy (anonymous):

48/53 gives the fraction of an hour ;)

OpenStudy (anonymous):

Distance = Speed * Time but we are looking for time so lets change the equation to : T = D/S D = 48 S = 53 plug in your info... T = 48/53 = .905 rounds to 0.9 or 9/10 of an hour